ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

20. Special Activities of the ICM 1998

1991 MS Classification: 68 Linton, Steve, University of St Andrews, United Kingdom GAP GAP is a free and open system for computational algebra and discrete mathematics, targetted primarily at researchers in these and related fields, but also used for research in other areas of mathematics and for teaching. The system provides fully integrated high quality implementations of stateof-the-art algorithms in a wide range of areas including group and semigroup theory, algebras especially Lie algebras, representation theory and combinatorics. A major feaure of the system is its extensibility. Much of the system is written in the GAP language, which is also available to users wishing to implement new algorithms, or add new types of mathematical object to the system. Such user contributed extensions which pass a refereeing process can be distributed with the system as share packages.